OpenStudy (anonymous):

A resuce ship spots a plane crash survivor 60m high on a cliff that is 250m (horizontally) from the ship. If they went to launch a small bundle with a radio and basic supplies, and the launcher is set at 40 degrees with respect to horizontal, how fast should they launch the bundle? Assume no air resistance.

OpenStudy (anonymous):

we have to hit the point x = 250m, y=60m y = v sin(40) t - 0.5 (9.8) t^2 x = v cos(40) t get time to cliff, t = 250/ v cos(40) put that t into y formula, set y = 60 and solve I got v = 59.0 m/s initial velocity t= 5.5s and hit the spot (within rounding errors)

OpenStudy (anonymous):

danke

OpenStudy (anonymous):

bitte...but check my calculations!

OpenStudy (anonymous):

well do auf Wiedersehen